# Auf der Morgenstelle 14, D-72076 Tübingen, Germany

### v1

@@ 1 Introduction
The most basic non-perturbative distribution functions describing the hadronic structure are parton distribution functions (PDFs) Owens1992 . They parametrize the probability of finding a parton of longitudinal momentum fraction $\xi$ in a parent hadron. Different PDFs exist for unpolarized Owens1992 , longitudinally polarized DeFlorian2008 , and transversely polarized hadrons Artru1990 . Their evolution equation is the so-called Dokshitzer-Gribov-Lipatov-Altarelli-Parisi (DGLAP) equation Dokshitzer1977 ; Dokshitzer1980 ; Gribov1972 ; Altarelli1977 , which will be discussed in more detail below.
@@
This paper is organized as follows. Section 2 summarizes basic notions of the DGLAP equation. In section 3 we introduce our semi-analytical $x$-space method and apply it to the non-singlet and singlet DGLAP equations, before briefly discussing the relation to Mellin-space. In section 4 we study the numerical performance of our method and compare the results to the Les Houches benchmark PDFs Alekhin:2002 as well as PDFs obtained using established Mellin-space methods. Section 5 concludes our paper and gives a brief outlook on ETQS evolution. Appendix A collects the master integrals needed for applying our method to the leading order (LO) DGLAP equations, while appendix C lists the input distributions we use for numerical evolution. This paper comes with Mathematica and Python implementations of the program POMPOM (Parton evolution with matrices: polynomials, logarithms, and more) which employ the presented method. They can be found as ancillary files on the arXiv.
@@ 2 DGLAP equation and evolution basis
In this section, we briefly introduce basic notions of the Dokshitzer-Gribov-Lipatov-Altarelli-Parisi (DGLAP) equation Dokshitzer1977 ; Dokshitzer1980 ; Gribov1972 ; Altarelli1977 . At LO, $f(x)$ is the probability density to find a parton $f$ carrying longitudinal momentum fraction $x$ of its parent hadron. This simple interpretation does not strictly hold at higher orders, where the parton densities become renormalization- and factorization-scheme dependent. Through factorization, the PDFs acquire a dependence on the scale $\mu$. How the PDFs change when the scale is varied, is described by the DGLAP equation. An intuitive derivation of the DGLAP equation can be found in standard textbooks Ellis1996 . More precise arguments based on the operator product expansion and the renormalization group Georgi1974 ; Gross1974 are given in Collins2011 .
@@
The next-to-leading order (NLO) evolution kernels can also be found in Ellis1996 . The complete DGLAP kernels are currently known to NNLO Vogt2004 ; Moch2004 , approximate results for the fourth-order (N3LO) quark-gluon splitting function $P_{gq}^{(3)}$ were published recently falcioni2024fourloop . The running coupling $\alpha_{s}$ is known to N3LO accuracy Deur2016 .
@@ 3.2 Solving the non-singlet and singlet DGLAP equation
$a_{MN}\!\left(\mu^{2}\right)=\,\exp\!\left[\Omega\!\left(\mu^{2},\mu_{0}^{2}\right)\right]a_{MN}\!\left(\mu_{0}^{2}\right),$ (26)
@@
We see that starting at NNLO, that is $\mathcal{O}(\alpha_{s}^{3})$, commutators of different $\mathcal{P}^{(n)}$ contribute. Expressions for $\Omega_{n\geq 3}$, which contain nested commutators of the $\mathcal{P}^{(n)}$, can be found in Blanes_2009 and references therein. Contributions due to $\Omega_{3}$ are at least of $\mathcal{O}(\alpha_{s}^{4})$.
@@
Note that we included $x^{-1}$ in our set of spanning functions to capture the known, more strongly divergent, small-$x$ behavior of the gluon and sea-quark PDF. Plugging this ansatz into eq. (10), we write the appearing integrals in terms of the master integrals $I_{i}$ defined in appendix A. As in the non-singlet case, we insert these master integrals and collect with respect to our spanning functions, which yields a differential equation for the coefficients,
@@ 4 Numerical analysis
The proposed $x$-space method to solve eqs. (10) is exact for an infinite number of basis functions. The key question is whether we can achieve sufficient accuracy for a realistic finite number of basis functions. The goal of this chapter is establishing the numerical validity of the presented method along with its implementation in POMPOM and to determine a finite set of basis functions giving optimal results. The criteria we will use to assess the quality of the evolution are a comparison with existing methods, consistency under back-evolution, violation of sum rules from truncation effects, and convergence when increasing the number of basis functions.
As discussed above, we need to select a finite set of basis functions to solve the infinite systems of coupled differential equations (24) and (30) numerically. As we will see, choosing different cut-offs has a sizeable effect. Hence, we start the numerical investigation of our method by determining a good cut-off $M_{\text{max}}$, $N_{\text{max}}(m)$ for the parametrization in eq. (32). Here, we use the Les Houches PDFs as a benchmark for the quality of our method Alekhin:2002 .
@@ 4.1 Choice of basis cut-off
To establish the numerical validity of the proposed method we compare with the Les Houches benchmark PDFs Alekhin:2002 . We observe that for a fixed total number of basis functions the quality of the agreement strongly depends on the choice of cut-off. For given $M_{\text{max}}$, we parametrize $N_{\text{max}}(m)$ as
@@
However, the optimal cut-off has to take into account not only the magnitude of the basis functions but also the magnitude of the coefficients. These are determined from the initial conditions and the respective elements of the evolution matrix. A typical evolution matrix for some choice of total number of basis functions and cut-off is depicted in figure 2.
@@ 4.2 Comparison to Les Houches benchmark tables
To determine the best cut-off, we optimize the parameters $\alpha$ and $\beta$ to minimize the deviation from the Les Houches benchmark tables Alekhin:2002 . These are commonly used for checking PDF evolution codes Candido2022 . Starting from $\alpha=0$, $\beta=0$ we perform a gradient descent on the lattice of integer values for $\alpha$ and $\beta$ until we reach a local minimum. This successive optimization of the cut-off is shown in figure 3.
@@ 4.3 Evolution of realistic PDF set
From the comparison with the Les Houches benchmark tables we have seen that POMPOM works properly and preserves sum rules to a high level of accuracy. To further showcase the capability of the proposed new method in the context of DGLAP evolution we change the initial conditions and turn to a more realistic model for PDFs. For the numerical evolution in this section, we use the leading order input distributions at $\mu_{0}^{2}=0.26\,\mathrm{GeV^{2}}$ from Gluck:1998xa , which are listed in appendix C and take the running coupling at leading order, which is given by eq. (73). Note that while these input distributions have an analytic parametrization in $x$, discrete input distributions can be interpolated with functions of the form (32) and subsequently evolved with the POMPOM method. Technical details of such an interpolation are beyond the scope of the present work. The evolution is performed in the variable flavor number scheme, where one evolves up to a flavor threshold, the resulting distribution is then taken as the new input above the threshold. At leading order there are no non-trivial matching conditions. The output distribution are immediately suitable to be new inputs without the need of additional steps. This is in contrast to an evolution in Mellin-space where evolution space and output space do not coincide and a transformation is required.
The PDFs from Gluck:1998xa provide us with sufficiently realistic initial conditions which are different from the benchmark to demonstrate that our method does not rely on specifics of the initial distributions. We note that due to the change of initial conditions, the optimal cut-off is also slightly different. As a way to fix the optimal cut-off in the absence of a previously known benchmark, we use the average deviation of flavor and momentum sum rules. As discussed in the comparison with the Les Houches benchmarks, these sum rules are a sensible measure for the global quality of the evolution.
@@ 4.4 Consistency of back-evolution
To perform a consistency check of our evolution code, we evolve the initial PDFs from the starting scale $\mu_{i}^{2}=0.26\,\mathrm{GeV^{2}}$ to $\mu^{2}=10^{8}\,\mathrm{GeV^{2}}$ and then back to $\mu_{i}$. Of course, for an exact evolution, the back-evolved PDF will exactly coincide with the initial PDF. However, any approximate method may result in errors of the order of the approximations applied.
@@ 4.5 Comparison to Mellin-space evolution
Here, we consider rather small numbers of basis functions. The idea is to investigate how many basis functions are required to obtain qualitatively plausible results. This is potentially interesting if one aims for lowering computational cost on the expense of accuracy, which might be desirable when considering the evolution of multi-parton distributions in the future. There, the basis needs to span multiple dimensions, resulting in a increased number of required functions.
Results for the comparison of POMPOM with different numbers of basis functions to Mellin evolution are given in figure 5.
@@
The evolution up to $\mu^{2}=10^{8}\,\mathrm{GeV^{2}}$ is performed in the variable flavor number scheme. The cut-off parameters are chosen as $\alpha=3$ and $\beta=4$, which are found to be the optimal numbers for $\sim 75$ basis functions.
@@ 4.6 Convergence for increasing number of basis functions
Figure 6: Absolute difference between singlet PDFs $x\,q_{\text{s}}(x)$ evolved from scale $\mu_{0}^{2}=0.4\,\mathrm{GeV^{2}}$ to $\mu^{2}=10^{8}\,\mathrm{GeV^{2}}$ with POMPOM for increasing number of basis functions. The cut-off parameters are $\alpha=\beta=2$.
Figure 7: Relative difference between singlet PDFs $x\,q_{\text{s}}(x)$ evolved from scale $\mu_{0}^{2}=0.4\,\mathrm{GeV^{2}}$ to $\mu^{2}=10^{8}\,\mathrm{GeV^{2}}$ with POMPOM for increasing number of basis functions. The cut-off parameters are $\alpha=\beta=2$.
In figures 6 and 7, respectively, we compare the absolute and relative differences for evolutions with between $78$ and $269$ basis functions in steps of $\sim 40$. We observe that the differences between steps becomes smaller and smaller indicating convergence towards zero. The absolute differences are largest for small $x$ since the singlet PDF grows strongly in this region. The lowest absolute differences occur at large $x\approx 1$. The relative differences are roughly constant for many orders of magnitude of $x$ between $10^{-7}$ and $10^{-1}$. For large $x\approx 1$ the relative difference deteriorates due to the smallness of the singlet PDF in this region. Overall, we observe satisfactory convergence for a modest linear increase in the number of basis functions. By adjusting the cut-off for higher numbers of basis functions, the rate of convergence could be further improved.

### v5

@@ 1 Introduction
The most basic non-perturbative distribution functions describing the hadronic structure are parton distribution functions (PDFs) Owens1992 . They parametrize the probability of finding a parton of longitudinal momentum fraction $\xi$ in a parent hadron. Different PDFs exist for unpolarized Owens1992 , longitudinally polarized DSSV2008 , and transversely polarized hadrons Artru1990 . Their evolution equation is the so-called Dokshitzer-Gribov-Lipatov-Altarelli-Parisi (DGLAP) equation Dokshitzer1977 ; Dokshitzer1980 ; Gribov1972 ; Altarelli1977 , which will be discussed in more detail below.
@@
This paper is organized as follows. Section 2 summarizes basic notions of the DGLAP equation. In section 3 we introduce our semi-analytical $x$-space method and apply it to the non-singlet and singlet DGLAP equations, before briefly discussing the relation to Mellin-space. In section 4 we study the numerical performance of our method and compare the results to the Les Houches benchmark PDFs Giele:2002hx as well as PDFs obtained using established Mellin-space methods. Section 5 concludes our paper and gives a brief outlook on ETQS evolution. Appendix A collects the master integrals needed for applying our method to the leading order (LO) DGLAP equations, appendix B gives the differential equation for the non-singlet coefficients explicitly, while appendix C lists the input distributions we use for numerical evolution. This paper comes with Mathematica and Python implementations of the program POMPOM (Parton evolution with matrices: polynomials, logarithms, and more) which employ the presented method. They can be found as ancillary files on the arXiv. Appendix D gives an estimate of the computation time of the code.
@@ 2 DGLAP equation and evolution basis
In this section, we briefly introduce basic notions of the DGLAP equation Dokshitzer1977 ; Dokshitzer1980 ; Gribov1972 ; Altarelli1977 . At LO, $f(x)$ is the probability density to find a parton $f$ carrying longitudinal momentum fraction $x$ of its parent hadron. This simple interpretation does not strictly hold at higher orders, where the parton densities become renormalization- and factorization-scheme dependent. Through factorization, the PDFs acquire a dependence on the scale $\mu$. How the PDFs change when the scale is varied, is described by the DGLAP equation. An intuitive derivation of the DGLAP equation can be found in standard textbooks Ellis1996 . More precise arguments based on the operator product expansion and the renormalization group Georgi1974 ; Gross1974 are given in Collins2011 .
@@
The next-to-leading order (NLO) evolution kernels can also be found in Ellis1996 . The complete DGLAP kernels are currently known to NNLO Vogt2004 ; Moch2004 ; Blumlein:2021enk , approximate results for the fourth-order (N3LO) quark-gluon splitting function $P_{gq}^{(3)}$ were published recently falcioni2024fourloop . The running coupling $\alpha_{s}$ is known to N4LO accuracy Baikov:2017 ; Herzog:2017 ; Luthe:2017 , a review on the empirical and theoretical knowledge of $\alpha_{s}$ can be found in Deur2016 .
@@ 3.2 Solving the non-singlet and singlet DGLAP equation
$a_{MN}\!\left(\mu^{2}\right)=\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\prod_{k=1}^{K}}\exp\!\left[\Omega\!\left(\mu^{2}_{k},\mu_{k-1}^{2}\right)\right]a_{MN}\!\left(\mu_{0}^{2}\right),\quad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{with }\mu^{2}_{K}=\mu^{2},}$ (26)
@@
We see that starting at NNLO, that is $\mathcal{O}(\alpha_{s}^{3})$, commutators of different $\mathcal{P}^{(n)}$ contribute. Expressions for $\Omega_{n\geq 3}$, which contain nested commutators of the $\mathcal{P}^{(n)}$, can be found in Blanes_2009 and references therein. Contributions due to $\Omega_{3}$ are at least of $\mathcal{O}(\alpha_{s}^{4})$. The intermediate points $\mu_{k}$ are required if $\mu$ is outside the domain of convergence of the Magnus expansion Blanes_2009 . They have to be chosen such that omitted terms in the expansion are sufficiently suppressed, in other words such that the counting of powers of $\alpha_{s}$ in the integrand is not spoiled by a large integration domain. The optimal number of slices $K$ will have to be determined on a case-by-case basis balancing between required accuracy and computation time. A detailed discussion of a Magnus expansion approach towards a differential equation of the form of eq. (21) can be found in Blanes_2009 , where its numerical performance is compared to standard Runge-Kutta methods. Of course the latter could also be used for solving eq. (24). We note that using the Magnus expansion in the context of the DGLAP equation has also been proposed recently in Simonelli:2024 to determine an analytical solution in Mellin-space beyond leading order.
@@
Note that we included $x^{-1}$ in our set of spanning functions to capture the known, more strongly divergent, small-$x$ behavior of the gluon and sea-quark PDFs. Plugging this ansatz into eq. (10), we write the appearing integrals in terms of the master integrals $I_{i}$ defined in appendix A. As in the non-singlet case, we insert these master integrals and collect with respect to our spanning functions, which yields a differential equation for the coefficients,
@@ 4 Numerical analysis
The proposed $x$-space method to solve eqs. (9) and (10) is exact for an infinite number of basis functions. The key question is whether we can achieve sufficient accuracy for a realistic finite number of basis functions. The goal of this chapter is to establish the numerical validity of the presented method along with its implementation in POMPOM and to determine a finite set of basis functions giving optimal results. The criteria we will use to assess the quality of the evolution are a comparison with existing methods, consistency under back-evolution, the magnitude of sum rule violations from truncation effects, and convergence when increasing the number of basis functions.
As discussed above, we need to select a finite set of basis functions to solve the infinite systems of coupled differential equations (24) and (30) numerically. As we will see, choosing different cut-offs has a sizeable effect. Hence, we start the numerical investigation of our method by determining a good cut-off $M_{\text{max}}$, $N_{\text{max}}(m)$ for the parametrization in eq. (32). Here, we use the Les Houches PDFs as a benchmark for the quality of our method Giele:2002hx .
@@ 4.1 Choice of basis cut-off
To establish the numerical validity of the proposed method we compare with the Les Houches benchmark PDFs Giele:2002hx . We observe that for a fixed total number of basis functions the quality of the agreement strongly depends on the choice of cut-off. For given $M_{\text{max}}$, we parametrize $N_{\text{max}}(m)$ as
@@
However, the optimal cut-off has to take into account not only the magnitude of the basis functions but also the magnitude of the coefficients. These are determined from the initial conditions and the respective elements of the evolution matrix. A typical evolution matrix for a particular choice of total number of basis functions and cut-off is depicted in figure 2.
@@ 4.2 Comparison to Les Houches benchmark tables
Figure 3: Comparison between POMPOM with 200 basis functions and the Les Houches benchmark table for an evolution from $\mu_{0}^{2}=2\,\mathrm{GeV^{2}}$ to $\mu_{\text{f}}^{2}=10^{4}\,\mathrm{GeV^{2}}$ in a fixed flavor number scheme with $n_{\mathrm{f}}=4$.
We start the investigation of the numerical importance of the basis cut-off by comparing POMPOM against the Les Houches benchmark tables Giele:2002hx . These are commonly used for checking PDF evolution codes Candido2022 . In figure 3 we show the results of the comparison between POMPOM with 200 basis functions and the Les Houches benchmark table for an evolution from $\mu_{0}^{2}=2\,\mathrm{GeV^{2}}$ to $\mu_{\text{f}}^{2}=10^{4}\,\mathrm{GeV^{2}}$ in a fixed flavor number scheme with $n_{\mathrm{f}}=4$. The POMPOM cut-off parameters are chosen as $(\alpha,\beta)=(2,11)$, this choice will be discussed in more detail below. In the figure on the right we observe that for $x\leq 0.5$ the relative difference is less than $10^{-4}$, for all $x\leq 0.3$ there is agreement in all significant digits. For the largest values of $x$ the relative deviations increase due to the smallness of the PDFs near near $x=1$. As the figure on the right demonstrates, the absolute error for $x\geq 0.5$ is below $10^{-6}$. Hence we see that even though the cut-off on basis functions in the POMPOM method may lead to large relative deviations in the vicinity of $x=1$, the absolute error is under control.
@@
To determine the best cut-off, we optimize the parameters $\alpha$ and $\beta$ to minimize the deviation from the Les Houches benchmark tables Giele:2002hx . Starting from the quadratic cut-off $(\alpha,\beta)=(0,0)$ we perform a gradient descent on the lattice of integer values for $\alpha$ and $\beta$ until we reach a local minimum. This successive optimization of the cut-off is shown in figure 4.
@@ 4.3 Evolution of realistic PDF set
From the comparison with the Les Houches benchmark tables we have seen that POMPOM works properly and preserves sum rules to a high level of accuracy. To further showcase the capability of the proposed new method in the context of DGLAP evolution we change the initial conditions and turn to a more realistic model for PDFs. For the numerical evolution in this section, we use the leading order input distributions at $\mu_{0}^{2}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0.26}\,\mathrm{GeV^{2}}$ from Gluck:1998xa , which are listed in appendix C and take the running coupling at leading order, which is given by eq. (73). Note that while these input distributions have an analytic parametrization in $x$, discrete input distributions can be interpolated with functions of the form (32) and subsequently evolved with the POMPOM method. Technical details of such an interpolation are beyond the scope of the present work. The evolution is performed in the variable flavor number scheme, where one evolves up to a flavor threshold, the resulting distribution is then taken as the new input above the threshold. At leading order there are no non-trivial matching conditions, which only come into play starting at NLO Bertone2014 ; Candido2022 . Here, the output distribution are immediately suitable to be new inputs without the need of additional steps. This is in contrast to an evolution in Mellin-space where evolution space and output space do not coincide and a transformation is required.
The PDFs from Gluck:1998xa provide us with sufficiently realistic initial conditions which are different from the benchmark to demonstrate that our method does not rely on specifics of the initial distributions. We note that due to the change of initial conditions, the optimal cut-off is also slightly different. As a way to fix the optimal cut-off in the absence of a previously known benchmark, we use the average violation of flavor and momentum sum rules. As discussed in the comparison with the Les Houches benchmarks, these sum rules are a sensible measure for the global quality of the evolution.
@@ 4.4 Consistency of back-evolution
Figure 6: Comparison of back-evolved PDF with initial condition. The red curve shows the difference between the exact initial condition $u_{v,\text{exact}}(x)$ and the initial condition expanded in a basis of 200 functions $u_{v,0}(x)$, which enters the evolution at $\mu_{0}^{2}=0.26\,\mathrm{GeV^{2}}$. The blue curve shows the difference between $u_{v,0}(x)$ and the back-evolved PDF $u_{v,\text{back}}(x)$, which has been evolved from $\mu_{0}^{2}=0.26\,\mathrm{GeV^{2}}$ to $\mu^{2}=10^{8}\,\mathrm{GeV^{2}}$ and back to $\mu_{0}^{2}=0.26\,\mathrm{GeV^{2}}$.
To perform a consistency check of our evolution code, we evolve the initial PDFs from the starting scale $\mu_{0}^{2}=0.26\,\mathrm{GeV^{2}}$ to $\mu^{2}=10^{8}\,\mathrm{GeV^{2}}$ and then back to $\mu_{0}$. Of course, for an exact evolution, the back-evolved PDF will exactly coincide with the initial PDF. However, any approximate method may result in errors of the order of the approximations applied. In figure 6 we show the difference between the initial condition expressed in the used basis $u_{v,0}(x)$ and the back-evolved PDF $u_{v,\text{back}}(x)$. Compared to the difference due to truncation between the exact initial condition $u_{v,\text{exact}}(x)$ and the basis representation $u_{v,0}(x)$ we can consider the back-evolution to be exact up to numerical inaccuracies. Absolute differences are plotted rather than relative ones to avoid division by very small values near $x=1$.
@@ 4.5 Comparison to Mellin-space evolution
First, we consider rather small numbers of basis functions. The idea is to investigate how many basis functions are required to obtain qualitatively plausible results. This is potentially interesting if one aims for lowering computational cost on the expense of accuracy, which might be desirable when considering the evolution of multi-parton distributions in the future. There, the basis needs to span multiple dimensions, resulting in a increased number of required functions. Results for the comparison of POMPOM with different numbers of basis functions to Mellin evolution are given in figure 7.
The evolution up to $\mu^{2}=10^{8}\,\mathrm{GeV^{2}}$ is performed in the variable flavor number scheme. The cut-off parameters are chosen as $\alpha=3$ and $\beta=4$, which are found to be the optimal numbers for $\sim 75$ basis functions. To showcase the qualitative features, absolute results are plotted on a linear scale. A more detailed look at the level of agreement is taken below.
@@
Figure 8: Relative and absolute differences between POMPOM and Mellin space evolution for non-singlet and singlet evolution. The left figure shows the relative differences on a logarithmic $x$-scale to visualize small-$x$ behavior, the right figure shows absolute differences on a linear $x$-scale to visualize large $x$ behavior.
Now we take a closer look at the level of agreement between POMPOM and Mellin evolution. For this purpose we take a moderately higher number of $200$ basis functions with cut-off parameters $(\alpha,\beta)=(2,2)$. Evolution is performed to $\mu^{2}=10^{8}\,\mathrm{GeV^{2}}$ as above, the relative and absolute differences between POMPOM and Mellin evolution are plotted in figure 8. To visualize the small-$x$ behavior, the left figure shows relative differences on a logarithmic scale. For the non-singlet evolution we find agreement below the $\sim 10^{-5}$ level for $10^{-7}\lesssim x\lesssim 0.2$ with minimal relative difference around $x=10^{-4}$. For the singlet evolution the agreement is on this level for $x\lesssim 0.1$ and is roughly constant down to $x=10^{-8}$. In the right figure, we look at the large-$x$ behavior on a linear $x$-scale. To avoid division by the small values of PDFs near $x=1$ absolute differences between POMPOM and Mellin Evolution are presented in this case. For large $x$ the absolute differences stay below $10^{-5}$ and do not increase when approaching $x=1$ confirming the findings from the comparison with the Les Houches benchmark in figure 3.
These comparisons show that POMPOM is in reasonable agreement with well-established methods for DGLAP evolution at leading order, demonstrating that the cut-off effect can be controlled with a numerically manageable number of basis functions. Since higher orders in $\alpha_{s}$ will presumably, due to their suppression in $\alpha_{s}$, only result in moderate corrections to the matrix elements used in POMPOM evolution, one can hope for a comparable level of numerical agreement for a certain number of basis functions. A detailed comparison to existing evolution codes such as HOPPET, APFEL, and EKO is beyond the scope of this paper since it will be meaningful only after POMPOM has been upgraded to a state-of-the-art DGLAP evolution code. Here we only aim for a proof-of-concept of the proposed method to justify its future developments for higher order DGLAP evolution or structurally similar evolution equations.
@@ 4.6 Convergence for increasing number of basis functions
Figure 9: Absolute difference between singlet PDFs $x\,q_{\text{s}}(x)$ evolved from scale $\mu_{0}^{2}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0.26}\,\mathrm{GeV^{2}}$ to $\mu^{2}=10^{8}\,\mathrm{GeV^{2}}$ with POMPOM for increasing number of basis functions. The cut-off parameters are $\alpha=\beta=2$.
In figures 9 and 10, respectively, we compare the absolute and relative differences for evolutions with between $78$ and $269$ basis functions in steps of $\sim 40$. We observe that the differences between steps becomes smaller and smaller indicating convergence towards zero. The absolute differences are largest for small $x$ since the singlet PDF grows strongly in this region. The lowest absolute differences occur at large $x\approx 1$. The relative differences are roughly constant for many orders of magnitude of $x$ between $10^{-7}$ and $10^{-1}$. For large $x\approx 1$ the relative difference deteriorates due to the smallness of the singlet PDF in this region. Overall, we observe satisfactory convergence for a modest linear increase in the number of basis functions. By adjusting the cut-off for higher numbers of basis functions, the rate of convergence could be further improved.
Figure 10: Relative difference between singlet PDFs $x\,q_{\text{s}}(x)$ evolved from scale $\mu_{0}^{2}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0.26}\,\mathrm{GeV^{2}}$ to $\mu^{2}=10^{8}\,\mathrm{GeV^{2}}$ with POMPOM for increasing number of basis functions. The cut-off parameters are $\alpha=\beta=2$.